Calculate phase shifts for quantum scattering, and then in terms of these phases give an account on scattering amplitude. As an application consider a sphere of radius r, assuming a certain potential V such that it is infinite inside the sphere and minimum outside the sphere,calculate phase shifts for this sphere.


1
Expert's answer
2020-07-03T08:28:48-0400

As per the question,

Radius of the sphere"=r"

Potential of the sphere"=V"

Phase shit- Angle "\delta_l" is known as the phase shift.

The partial wave equation of the plane wave "e^{ikt}=\Sigma_{l=0}^\infty\sqrt{4\pi(l+1)}i^lj_l(K_r)Y_{lo}(\theta)"


We know that asymptotically spherical Bessel function, "j_l{(kr)}" has the form,

"j_l{(kr)} \ \ \overrightarrow{r\rightarrow\infty} \ \ \frac{\sin(kr-\frac{l\pi}{2})}{kr}"


Now, substituting the value and solving it further,

"e^{ikt}=\Sigma_{l=0}^\infty\sqrt{4\pi(l+1)}i^lj_l(K_r)Y_{lo}(\theta)"


"\Rightarrow e^{ikt}=\Sigma_{l=0}^\infty\sqrt{4\pi(l+1)}\{\frac{e^{ikr}}{2ikr}-\frac{i^le^{-i(kr-\frac{l\pi}{2}))}}{2ikr} \}Y_{lo}(\theta)"